# Artificial Precision Polarization Array: Sensitivity for the axion-like dark matter with clock satellites

### v1

@@ Appendix B GLSP: Generalized Lomb-Scargle Periodogram
Under the APPA framework, the satellite network is confined to a small scale region of the Solar System with $L_{i}\sim\mathcal{O}\big(10^{5}\big)\mathrm{km}$. For axions velocity $v\sim 10^{-3}c$, the de Broglie wavelength $\lambda_{dB}$ of the scalar axion field is $\lambda_{de}=\frac{h}{m_{a}v}\sim\mathcal{O}\big(10^{-8}-10^{3}\big)\mathrm{pc}$, which is much larger than the spatial extent of the satellite array for axion masses in the range $m_{a}\sim\mathcal{O}\big(10^{-22}-10^{-12}\big)$eV. Therefore, the entire network can be treated as residing within a single coherent axion field, implying $\varphi_{o}=\varphi_{s}=\varphi$ and $\rho_{o}=\rho_{s}=\rho_{\text{DM}}$. Consequently, a significant fraction of environmental disturbances and intrinsic signal uncertainties are effectively eliminated, and Eq. (II.10) reduces to
@@
Since the satellite network is distributed on a spatial scale of $10^{5}$ $\mathrm{km}$, the corresponding signal transmission time is $T\sim\mathcal{O}\big(1\big)$s. Thus, for the axion masses $m_{a}\leq\mathcal{O}\big(10^{-16}\big)$eV, one has $\mathrm{cos}\big(m_{a}T\big)\simeq 1$. Under this condition, Eq. (B.1) can be rewritten as
@@
For the axion mass $m_{a}\sim\mathcal{O}\big(10^{-16}-10^{-12}\big)$eV, Eq. (B.2) no longer applies, the full expression $\big[\left(\alpha_{o}^{2}+\alpha_{s}^{2}\right)-2\alpha_{o}\alpha_{s}\mathrm{cos}\left(m_{a}T\right)\big]^{\frac{1}{2}}$ must be considered. The corresponding distribution results are shown in Fig. 8. Note that for arbitrary axion mass, the statistical distribution of amplitude fluctuations lies between the two limiting cases of $\mathrm{cos}\big(m_{a}T\big)=1$ (red reigon in Fig. 8) and $\mathrm{cos}\big(m_{a}T\big)=-1$ (purple reigon in Fig. 8).
Figure 8: The Histograms and corresponding PDFs of $\big[\left(\alpha_{o}^{2}+\alpha_{s}^{2}\right)-2\alpha_{o}\alpha_{s}\mathrm{cos}\left(m_{a}T\right)\big]^{\frac{1}{2}}$ for different axion masses $m_{a}$. The distributions were generated using the MC method with $10^{6}$ samples for each $m_{a}$. Color coding: red, $\mathrm{cos}\big(m_{a}T\big)=1$; Orange, $m_{a}=10^{-15}\mathrm{eV}$; Yellow, $m_{a}=10^{-14}\mathrm{eV}$; Green, $m_{a}=10^{-13}\mathrm{eV}$; Blue, $m_{a}=10^{-12}\mathrm{eV}$; Purple, $\mathrm{cos}\big(m_{a}T\big)=-1$. The black curve represents the envelope defined by the maxima of the PDFs across different values of $m_{a}$.
@@
$P_{ab}$ $=\overline{\Delta\phi_{a}(t)\Delta\phi_{b}(t)}$ (B.3)
$=\left(\frac{g_{a\gamma}}{m_{a}}\sqrt{\frac{\rho_{\text{DM}}}{2}}\right)^{2}\Big\langle\left|\alpha_{o}-\alpha_{sa}\right|\cos\left(m_{a}t+\varphi_{a}\right)$
$\quad\quad\quad\quad\qquad\left|\alpha_{o}-\alpha_{sb}\right|\cos\left(m_{a}t+\varphi_{b}\right)\Big\rangle$
$\propto\Big\langle\left|\alpha_{o}-\alpha_{sa}\right|\left|\alpha_{o}-\alpha_{sb}\right|$
$\quad\frac{1}{2}\big[\cos\big(2m_{a}t+\varphi_{a}+\varphi_{b}\big)+\cos\big(\varphi_{a}-\varphi_{b}\big)\big]\Big\rangle$
$=\frac{1}{2}\big\langle\left|\alpha_{o}-\alpha_{sa}\right|\left|\alpha_{o}-\alpha_{sb}\right|\cos\big(\varphi_{a}-\varphi_{b}\big)\big\rangle.$
@@
Figure 9: The histogram of Eq. (B.3) (yellow, 100 bins) together with its PDF (blue) is shown. A total of $10^{6}$ samples were generated using the MC method.

### v2

@@ Appendix B GLSP: Generalized Lomb-Scargle Periodogram
Under the APPA framework, the satellite network is confined to a small scale region of the Solar System with $L_{i}\approx 5$AU (Jupiter’s orbit). For axions velocity $v\sim 10^{-3}c$, the de Broglie wavelength $\lambda_{\text{de}}$ of the scalar axion field is $\lambda_{\text{de}}=\frac{h}{m_{a}v}\sim\mathcal{O}\big(10^{-5}-10^{2}\big)\mathrm{pc}$, which is much larger than the spatial extent of the satellite array for axion masses in the range $m_{a}\sim\mathcal{O}\big(10^{-22}-10^{-15}\big)$eV. Therefore, the entire network can be treated as residing within a single coherent axion field, implying $\varphi_{o}=\varphi_{s}=\varphi$ and $\rho_{o}=\rho_{s}=\rho_{\text{DM}}$. Consequently, a significant fraction of environmental disturbances and intrinsic signal uncertainties are effectively eliminated, and Eq. (II.10) reduces to
@@
Since the satellite network is distributed on a spatial scale of 55AU, the corresponding signal transmission time is $T\sim\mathcal{O}\big(10^{3}\big)$s. Thus, for the axion masses $m_{a}\leq 2.6\times 10^{-20}$eV, one has $\mathrm{cos}\big(m_{a}T\big)\simeq 1$. Under this condition, Eq. (B.1) can be rewritten as
@@
Figure 7: The Histograms and corresponding PDFs of $\big[\left(\alpha_{o}^{2}+\alpha_{s}^{2}\right)-2\alpha_{o}\alpha_{s}\mathrm{cos}\left(m_{a}T\right)\big]^{\frac{1}{2}}$ for different axion masses $m_{a}$. The distributions were generated using the MC method with $10^{6}$ samples for each $m_{a}$. Color coding: Red, $\mathrm{cos}\big(m_{a}T\big)=1$; Orange, $m_{a}=10^{-19}\mathrm{eV}$; Yellow, $m_{a}=10^{-18}\mathrm{eV}$; Green, $m_{a}=10^{-17}\mathrm{eV}$; Blue, $m_{a}=10^{-16}\mathrm{eV}$; Purple, $\mathrm{cos}\big(m_{a}T\big)=-1$. The black curve represents the envelope defined by the maxima of the PDFs across different values of $m_{a}$.
For the axion mass $m_{a}\geq 2.6\times 10^{-20}$eV, Eq. (B.2) no longer applies, the full expression $\big[\left(\alpha_{o}^{2}+\alpha_{s}^{2}\right)-2\alpha_{o}\alpha_{s}\mathrm{cos}\left(m_{a}T\right)\big]^{\frac{1}{2}}$ must be considered. The corresponding distribution results are shown in Fig. 7. Note that for arbitrary axion mass, the statistical distribution of amplitude fluctuations lies between the two limiting cases of $\mathrm{cos}\big(m_{a}T\big)=1$ (red reigon in Fig. 7) and $\mathrm{cos}\big(m_{a}T\big)=-1$ (purple reigon in Fig. 7).
@@
$P_{ab}$ $=\overline{\Delta\phi_{a}(t)\Delta\phi_{b}(t)}=\left(\frac{g_{a\gamma}}{m_{a}}\sqrt{\frac{\rho_{\text{DM}}}{2}}\right)^{2}\Big\langle\left|\alpha_{o}-\alpha_{sa}\right|$
$\quad\times\cos\left(m_{a}t+\varphi_{a}\right)\left|\alpha_{o}-\alpha_{sb}\right|\cos\left(m_{a}t+\varphi_{b}\right)\Big\rangle$
$\propto\frac{1}{2}\big\langle\left|\alpha_{o}-\alpha_{sa}\right|\left|\alpha_{o}-\alpha_{sb}\right|\cos\big(\varphi_{a}-\varphi_{b}\big)\big\rangle.$ (B.3)
@@
Figure 8: The histogram of Eq. (B) (yellow, 100 bins) together with its PDF (blue) is shown. A total of $10^{6}$ samples were generated using the MC method.